OpenStudy (anonymous):

The angles shown below are supplementary: A pair of supplementary angles is shown. One angle measures 10x and the other angle measures 60 degrees. What is the value of x? 6° 9° 12° 18°

OpenStudy (anonymous):

I wanted someone to check my answer which was 6

OpenStudy (anonymous):

add the angles and equate it to 180 10x + 60 = 180 .... 10x + 60 = 180 - 60 -60 -------------- 10x + 0 = 120 so, now we can focus on: 10x = 120 the 10x = 120, when x=12 So C

OpenStudy (anonymous):

okay thanks @_greatmath7

OpenStudy (anonymous):

yeah of course

OpenStudy (anonymous):

:)